Evaluate ∫{1/logx – 1/(logx)2}dx..Answer plz ..be quick i have my exams tomorrow

Grade:12

1 Answers

Shashank Singh
6 years ago
put logx=t
then x= e^t 1/x dx = dt
dx= e^t dt
now the expression becomes
e^t{f(t) + f’(t)} using this identity you will get the answer